# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

### v1

@@ 3.1 Problem formulation and solution
As noted in Remark 3.3, we have only characterised the solutions of (10) that satisfy (12) on $\operatorname{dom}{C}^{\dagger}$. That is, we have characterised $\mathcal{S}_{r,*}\cap\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\}$. We can now charachterise all solutions of (10), that need not satisfy the minimality property (12), on $\operatorname{dom}{C}^{\dagger}$.
@@
$\mathcal{S}_{r}\cap\left\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\right\}=\left\{\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}}:\ \hat{X}\in\mathcal{S}_{r,*},\right.$
$\left.T:\operatorname{dom}{C}^{\dagger}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},S:\operatorname{ran}{C}^{\perp}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}\right\}.$
@@
It follows directly from Theorem 3.2 that (10) has a solution if and only if $\mathcal{Y}_{r}\not=\emptyset$. Now let $X=\hat{X}+SP_{\operatorname{ran}{C}^{\perp}}+P_{\ker{B}}T$ with $\hat{X}$, $S$ and $T$ as specified. Then $\operatorname{dom}{X}=\operatorname{dom}{\hat{X}}=\operatorname{dom}{C}^{\dagger}$ and $BXC=B\hat{X}C$. But $\hat{X}\in\mathcal{S}_{r}$, so $X\in\mathcal{S}_{r}$. This shows the $\supset$” inclusion holds. Conversely, let $X\in\mathcal{S}_{r}$ with $\operatorname{dom}{X}=\operatorname{dom}{C^{\dagger}}$. We define $\hat{X}=B^{\dagger}BXCC^{\dagger}$, so $\hat{X}\in\mathcal{S}_{r,*}$ by Lemma 3.7 Items (ii) and (iii). In particular, using also (36) and (37), we have on $\operatorname{dom}{C}^{\dagger}=\operatorname{dom}{X}$ that $P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}=B^{\dagger}BXCC^{\dagger}=\hat{X}$. We also define $S\coloneqq XP_{\operatorname{ran}{C}^{\perp}}$ and $T\coloneqq{\left.\kern-1.2ptP_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{C}^{\dagger}}}$. So $S$ and $T$ well-defined, as $P_{\operatorname{ran}{C}^{\perp}}$ and ${\left.\kern-1.2ptP_{\overline{\operatorname{ran}{C}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{C}^{\dagger}}}$ both map into $\operatorname{dom}{C}^{\dagger}=\operatorname{dom}{X}$. Furthermore, $S=SP_{\operatorname{ran}{C}^{\perp}}$ and $T=P_{\ker{B}}T$. Hence using (17):
@@
Let $\tilde{X}$ be a solution of (10). By Corollary 3.9, we have that $\mathcal{S}_{r,*}\not=\emptyset$ and on $\operatorname{dom}{C^{\dagger}}$ that $\tilde{X}=\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}}$ for some $\hat{X}\in\mathcal{S}_{r,*}$, $T:\operatorname{dom}{C}^{\dagger}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}$ and $S:\operatorname{ran}{C}^{\perp}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}$. We have $\hat{X}=P_{\ker{B}^{\perp}}\hat{X}$ by (12). For $h\in\operatorname{ran}{C}$ it holds that $P_{\operatorname{ran}{C}^{\perp}}h=0$, hence,
@@
Let us assume $\hat{X}$ is unbounded. Then there exists a bounded sequence $(h_{n})_{n}$ in $\operatorname{dom}{C^{\dagger}}$ satisfying $\lVert\hat{X}h_{n}\rVert\rightarrow\infty$. Since $\hat{X}$ satisfies (12), $\hat{X}h_{n}=\hat{X}P_{\operatorname{ran}{C}}h_{n}$. So $k_{n}\coloneqq P_{\operatorname{ran}{C}}h_{n}\in\operatorname{ran}{C}$ is a bounded sequence and $\lVert\hat{X}k_{n}\rVert$ diverges. But then also $\lVert\tilde{X}k_{n}\rVert$ diverges, showing that $\tilde{X}$ is unbounded. ∎
@@ 3.2 The case of no solutions and approximate minimisers
Even when the condition $\mathcal{Y}_{r}\not=\emptyset$ with $\mathcal{Y}_{r}$ defined in 3.1 fails to hold, that is, even when $\operatorname{ran}{Y}\subset\operatorname{dom}{B}^{\dagger}$ does not hold for any rank-$r$ reduced SVD $Y$ of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$, one can still construct a minimizing sequence for (10) that satisfies (12). To do so, suppose $Y\coloneqq(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ has SVD $\sum_{i=1}^{r}\lambda_{i}f_{i}\otimes e_{i}$, with $(e_{i})_{i=1}^{n}$ and $(f_{i})_{i=1}^{n}$ orthonormal sequences in $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$ respectively. Since $f_{i}\in\operatorname{ran}{Y}\subset\overline{\operatorname{ran}{B}}$, we can find $f_{i}^{m}\in\overline{\operatorname{ran}{B}}$ such that $\lVert f_{i}-f_{i}^{m}\rVert<1/n$ whenever $m\geq n$ and $n\in\mathbb{N}$. Let us define $Y^{(n)}\coloneqq\sum_{i=1}^{r}\lambda_{i}f_{i}^{n}\otimes e_{i}$. Then,
$\lVert Y-Y^{(n)}\rVert_{HS}^{2}=\sum_{i=1}^{r}\lambda_{i}\lVert f_{i}^{n}-f_{i}\rVert^{2}\leq r\lambda_{1}n^{-2}\rightarrow 0,\quad n\rightarrow\infty.$
@@
$\inf\{\lVert M-BXC\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}\}$ $\leq\lim_{n}\lVert M-B\hat{X}^{(n)}C\rVert_{HS}^{2}$
$=\lim_{n}\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\hat{X}^{(n)}C\rVert_{HS}^{2}+c$
$=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{HS}^{2}+c$
$\leq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{HS}^{2}+c$
$=\lVert M-BXC\rVert_{HS}^{2}.$
By taking the infimum over $X\in\mathcal{X}_{r}$ we see that $\lVert M-B\hat{X}^{(n)}C\rVert_{HS}$ converges to $\inf\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{X}_{r}\}$.
@@ 3.3 Optimal approximation error
The quality of the optimal approximation (11) is addressed next.
@@
$\lVert M-B\hat{X}C\rVert_{HS}^{2}=\lVert M\rVert_{HS}^{2}-\Delta,$
with $\Delta\coloneqq\sum_{i\leq r}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}$. We have,
$\Delta=\sum_{i\leq r}\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MC^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}})=\sum_{i\leq r}\sigma_{i}(C^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}}MC^{\dagger}C)$
If additionally $\operatorname{ran}{M}\subset\operatorname{dom}{B^{\dagger}}$, then we may write $\Delta=\sum_{i\leq r}\lambda_{i}(B^{\dagger}MC^{\dagger}CM^{*}B),$ where $\lambda_{i}(\cdot)$ denotes taking the $i$-th largest eigenvalue.
@@
By Lemma 3.7 item (ii), $B\hat{X}C$ is a rank-$r$ reduced SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. So, by Lemma 3.6,
$\lVert M-BXC\rVert_{HS}^{2}$ $=\lVert M\rVert_{HS}^{2}+\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\hat{X}C\rVert_{HS}^{2}-\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\rVert_{HS}^{2}$
$=\lVert M\rVert_{HS}^{2}+\sum_{i>r}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}-\sum_{i}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}$
$=\lVert M\rVert_{HS}^{2}-\Delta.$
By Proposition A.8, and the fact that $P_{\overline{\operatorname{ran}{B}}}$ and $P_{\ker{C}^{\perp}}$ are orthogonal projectors,
@@
and $T\in L_{2}(\mathcal{H}_{2})$, $T=T^{*}$, $T\geq 0$. In particular, the singular values of $T$ equal the eigenvalues of $T$. To compute $\Delta$ under the assumption $\operatorname{ran}{M}\subset\operatorname{dom}{B}^{\dagger}$, we may thus use the eigenvalues of $T$. Suppose that there exists nonzero $v$ such that $Tv=\lambda v$, so $v\in\operatorname{ran}{B}\subset\operatorname{dom}{B^{\dagger}}$. With $w\coloneqq B^{\dagger}v$ it holds that $w\not=0$ and by (35) and (37), $B^{\dagger}MC^{\dagger}CM^{*}Bw=B^{\dagger}Tv=\lambda w$. So $\lambda$ is an eigenvalue of $B^{\dagger}MC^{\dagger}CM^{*}B$. Conversely, suppose there exists nonzero $w$ such that $B^{\dagger}MC^{\dagger}CM^{*}Bw=\lambda w$, so $w\not\in\ker{B}$. With $v\coloneqq Bw$ it holds that $v\not=0$ and by (34) and (37), $Tv=BB^{\dagger}MC^{\dagger}CM^{*}Bw=B\lambda w=\lambda v$. So $\lambda$ is an eigenvalue of $T$. We conclude that
@@ 4 Unbounded solution
Let $\mathcal{H}$ be a separable Hilbert space and take $\mathcal{H}_{i}=\mathcal{H}$. Let $(e_{n})_{n}$ be an orthonormal basis of $\mathcal{H}$. Let $\0(\gamma)_{n},(\mu_{n})_{n}\in\mathbb{C}\backslash\{0\}$ converge to zero. We assume $(\mu_{n})_{n}$ is nonincreasing. Let $(\alpha_{n})_{n}\in\mathbb{C}$ such that $\lvert\alpha_{n}\rvert\rightarrow\infty$ and $(\alpha_{n}\gamma_{n})_{n}$ is square summable. For example, $\gamma_{n}=n^{-2}$ and $\alpha_{n}=n$ satisfy these conditions.
We may then define $w\coloneqq\sum_{n\geq 2}\alpha_{n}\gamma_{n}e_{n}\in\mathcal{H}$. Put $f_{1}\coloneqq w\lVert w\rVert^{-1}$ and $f_{2}\coloneqq e_{1}$ and extend $\{f_{1},f_{2}\}$ to an ONB $\{f_{n}\}_{n}$ of $\mathcal{H}$. We then set, for $h\in\mathcal{H}$,
@@
Then $C$ is injective and $\operatorname{ran}{C}=\{h\in\mathcal{H}:\ \sum_{n}\gamma_{n}^{2}\langle h,e_{n}\rangle^{2}<\infty\}$. As $\operatorname{ran}{C}$ contains $\{e_{n}\}_{n}$, it is dense in $\mathcal{H}$, thus $\operatorname{ran}{C}^{\perp}=\{0\}$. Therefore, $\operatorname{dom}{C^{\dagger}}=\operatorname{ran}{C}$ and the Moore–Penrose inverse of $C$ is given, as in Example A.6, to be
@@
Let us take $B=I$ for simplicity, so $P_{\overline{\operatorname{ran}{B}}}=I$. As $C$ is injective, $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}=M$. Since $(\mu_{n})_{n}$ is nonincreasing, for any $r\geq 1$, one of the rank-$r$ truncated SVDs of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ is then given by,
@@
$\geq\big{\lvert}\mu_{1}\gamma^{-1}_{m}\langle e_{m},f_{1}\rangle\big{\rvert}=\lvert\mu_{1}\rvert\lvert\alpha_{m}\rvert\lVert w\rVert^{-1}\rightarrow\infty$
as $m\rightarrow\infty$, showing that $B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger}$ is unbounded. This shows that (11) can indeed give unbounded solutions. If $\mu_{1}>\mu_{2}$ and $r=1$, then (11) gives a unique solution, which shows that (11) can also give only unbounded solutions. By Corollary 3.10 it follows that in this case in fact all solutions of (10) are unbounded.
Now let $r=1$ and $\mu_{1}=\mu_{2}$, so that $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ has at least two distinct rank-$1$ truncated SVDs. The rank-$1$ truncated SVD (18), i.e. $(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{1}=\mu_{1}\langle\cdot,f_{1}\rangle f_{1}$, yields the following solution, which is of the form (11),
@@
We have thus constructed two solutions of the form (11), yet one of these is bounded, in particular it has finite rank and has finite Hilbert–Schmidt norm, and one is unbounded.

### v3

@@ 3.1 Problem formulation and solution
As noted in Remark 3.3, we have characterised $\mathcal{S}_{r,*}\cap\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\}$. We can now characterise $\mathcal{S}_{r}\cap\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\}$, that is, we characterise all solutions of (10) on $\operatorname{dom}{C^{\dagger}}$, that need not satisfy the minimality property (12).
@@
$\mathcal{S}_{r}\cap\left\{\operatorname{dom}{X}=\operatorname{dom}{C}^{\dagger}\right\}$
$=$ $\left\{\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}}:\ \hat{X}\in\mathcal{S}_{r,*},\ T:\operatorname{dom}{C}^{\dagger}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3},\ S:\operatorname{ran}{C}^{\perp}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}\right\}.$
@@
It follows directly from Theorem 3.2 that (10) has a solution if and only if $\mathcal{Y}_{r}\not=\emptyset$. Now let $X=\hat{X}+SP_{\operatorname{ran}{C}^{\perp}}+P_{\ker{B}}T$ with $\hat{X}$, $S$ and $T$ as specified. Then $\operatorname{dom}{X}=\operatorname{dom}{\hat{X}}=\operatorname{dom}{C}^{\dagger}$ and $BXC=B\hat{X}C$. But $\hat{X}\in\mathcal{S}_{r}$, so $X\in\mathcal{S}_{r}$. This shows the $\supset$” inclusion holds. Conversely, let $X\in\mathcal{S}_{r}$ with $\operatorname{dom}{X}=\operatorname{dom}{C^{\dagger}}$. We define $\hat{X}=B^{\dagger}BXCC^{\dagger}$, so $\hat{X}\in\mathcal{S}_{r,*}$ by Lemma 3.7 Items (ii) and (iii). In particular, using also (38) and (39), we have on $\operatorname{dom}{C}^{\dagger}=\operatorname{dom}{X}$ that $P_{\ker{B}^{\perp}}XP_{\overline{\operatorname{ran}{C}}}=B^{\dagger}BXCC^{\dagger}=\hat{X}$. We also define $S\coloneqq XP_{\operatorname{ran}{C}^{\perp}}$ and $T\coloneqq{\left.\kern-1.2ptP_{\ker{B}}XP_{\overline{\operatorname{ran}{C}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{C}^{\dagger}}}$. Thus, $S$ and $T$ are well-defined, as $P_{\operatorname{ran}{C}^{\perp}}$ and ${\left.\kern-1.2ptP_{\overline{\operatorname{ran}{C}}}\vphantom{\big{|}}\right|_{\operatorname{dom}{C}^{\dagger}}}$ both map into $\operatorname{dom}{C}^{\dagger}=\operatorname{dom}{X}$. Furthermore, $S=SP_{\operatorname{ran}{C}^{\perp}}$ and $T=P_{\ker{B}}T$. Hence using (17):
@@
Let $\tilde{X}$ be a solution of (10). By Theorem 3.2 we have $\mathcal{Y}_{r}\not=\emptyset$. Thus, by Lemma 3.7(iii), $\mathcal{S}_{r,*}\not=\emptyset$. By Corollary 3.9, it then holds on $\operatorname{dom}{C^{\dagger}}$ that $\tilde{X}=\hat{X}+P_{\ker{B}}T+SP_{\operatorname{ran}{C}^{\perp}}$ for some $\hat{X}\in\mathcal{S}_{r,*}$, $T:\operatorname{dom}{C}^{\dagger}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}$ and $S:\operatorname{ran}{C}^{\perp}\subset\mathcal{H}_{2}\rightarrow\mathcal{H}_{3}$. We have $\hat{X}=P_{\ker{B}^{\perp}}\hat{X}$ by (12). For $h\in\operatorname{ran}{C}$ it holds that $P_{\operatorname{ran}{C}^{\perp}}h=0$, hence,
@@
Let us assume $\hat{X}$ is unbounded. Then there exists a bounded sequence $(h_{n})_{n}$ in $\operatorname{dom}{C^{\dagger}}$ satisfying $\lVert\hat{X}h_{n}\rVert\rightarrow\infty$. Since $\hat{X}$ satisfies (12), $\hat{X}h_{n}=\hat{X}P_{\operatorname{ran}{C}}h_{n}$. Hence $k_{n}\coloneqq P_{\operatorname{ran}{C}}h_{n}\in\operatorname{ran}{C}$ is a bounded sequence and $\lVert\hat{X}k_{n}\rVert$ diverges. But then also $\lVert\tilde{X}k_{n}\rVert$ diverges, showing that $\tilde{X}$ is unbounded. ∎
@@ 3.2 The case of no solutions and approximate minimisers
Even when the condition $\mathcal{Y}_{r}\not=\emptyset$ with $\mathcal{Y}_{r}$ defined in 3.1 fails to hold, that is, even when $\operatorname{ran}{Y}\subset\operatorname{dom}{B}^{\dagger}$ does not hold for any rank-$r$ reduced SVD $Y$ of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$, one can still construct a minimizing sequence for (10) that satisfies (12). To do so, suppose $Y\coloneqq(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ has SVD $\sum_{i=1}^{r}\lambda_{i}f_{i}\otimes e_{i}$, with orthonormal sequences $(e_{i})_{i=1}^{r}$ and $(f_{i})_{i=1}^{r}$ in $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$ respectively. We recall the notation $\otimes$’ from Appendix A. Since $f_{i}\in\operatorname{ran}{Y}\subset\overline{\operatorname{ran}{B}}$, we can find $f_{i}^{m}\in{\operatorname{ran}{B}}$ such that $\lVert f_{i}-f_{i}^{m}\rVert<1/n$ whenever $m\geq n$ and $n\in\mathbb{N}$. Let us define $Y^{(n)}\coloneqq\sum_{i=1}^{r}\lambda_{i}f_{i}^{n}\otimes e_{i}$. Then,
$\lVert Y-Y^{(n)}\rVert_{L_{2}}^{2}=\sum_{i=1}^{r}\lambda_{i}^{2}\lVert f_{i}^{n}-f_{i}\rVert^{2}\leq r\lambda_{1}^{2}n^{-2}\rightarrow 0,\quad n\rightarrow\infty.$
@@
$\inf\{\lVert M-BXC\rVert_{L_{2}}^{2}:\ X\in\mathcal{X}_{r}\}$ $\leq\lim_{n}\lVert M-B\hat{X}^{(n)}C\rVert_{L_{2}}^{2}$
$=\lim_{n}\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\hat{X}^{(n)}C\rVert_{L_{2}}^{2}+c$
$=\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-Y\rVert_{L_{2}}^{2}+c$
$\leq\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-BXC\rVert_{L_{2}}^{2}+c$
$=\lVert M-BXC\rVert_{L_{2}}^{2}.$
By taking the infimum over $X\in\mathcal{X}_{r}$ we see that $\lVert M-B\hat{X}^{(n)}C\rVert_{L_{2}}$ converges to $\inf\{\lVert M-BXC\rVert_{L_{2}}:\ X\in\mathcal{X}_{r}\}$. We conclude that $(\hat{X}^{(n)})_{n}$ is a sequence of approximate minimisers, i.e. $\lVert M-B\hat{X}^{(n)}C\rVert_{L_{2}}\rightarrow\inf\{\lVert M-BXC\rVert_{L_{2}},\ X\in\mathcal{X}_{r}\}$, and each element in this sequence satisfies the minimality property $\hat{X}^{(n)}=P_{\ker{B}^{\perp}}\hat{X}^{(n)}P_{\overline{\operatorname{ran}{C}}}$ on $\operatorname{dom}{\hat{X}^{(n)}}=\operatorname{dom}{C}^{\dagger}.$
@@ 3.3 Optimal approximation error
The quality of the optimal approximation (11) is addressed in the following result. The result also implies that the optimal approximation error converges to 0 as $r\rightarrow\infty$.
@@
$\lVert M-B\hat{X}C\rVert_{L_{2}}^{2}=\lVert M\rVert_{L_{2}}^{2}-\Delta,$
with $\Delta\coloneqq\sum_{i\leq r}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}$. We have
$\Delta=\sum_{i\leq r}\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MC^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}})=\sum_{i\leq r}\sigma_{i}(C^{\dagger}CM^{*}P_{\overline{\operatorname{ran}{B}}}MC^{\dagger}C).$
If additionally $\operatorname{ran}{M}\subset\operatorname{dom}{B^{\dagger}}$, then $B^{\dagger}MC^{\dagger}CM^{*}B$ has only nonnegative eigenvalues and $\Delta=\sum_{i\leq r}\lambda_{i}(B^{\dagger}MC^{\dagger}CM^{*}B),$ where $\lambda_{i}(\cdot)$ denotes the $i$-th largest eigenvalue.
@@
By Lemma 3.7 item (ii), $B\hat{X}C$ is a rank-$r$ reduced SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. Thus, by Lemma 3.6,
$\lVert M-B\hat{X}C\rVert_{L_{2}}^{2}$ $=\lVert M\rVert_{L_{2}}^{2}+\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}-B\hat{X}C\rVert_{L_{2}}^{2}-\lVert P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}\rVert_{L_{2}}^{2}$
$=\lVert M\rVert_{L_{2}}^{2}+\sum_{i>r}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}-\sum_{i}\lvert\sigma_{i}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})\rvert^{2}$
$=\lVert M\rVert_{L_{2}}^{2}-\Delta.$
By Proposition A.8 and the fact that $P_{\overline{\operatorname{ran}{B}}}$ and $P_{\ker{C}^{\perp}}$ are orthogonal projectors,
@@
and $T\in L_{2}(\mathcal{H}_{2})$, $T=T^{*}$, $T\geq 0$. In particular, the singular values of $T$ equal the eigenvalues of $T$. To compute $\Delta$ under the assumption $\operatorname{ran}{M}\subset\operatorname{dom}{B}^{\dagger}$, we may thus use the eigenvalues of $T$. Suppose that there exists some nonzero $v$ such that $Tv=\lambda v$. Then $v\in\operatorname{ran}{B}\subset\operatorname{dom}{B^{\dagger}}$. With $w\coloneqq B^{\dagger}v$ it holds that $w\not=0$. By (37) and (39), $B^{\dagger}MC^{\dagger}CM^{*}Bw=B^{\dagger}Tv=\lambda w$. Thus $\lambda$ is an eigenvalue of $B^{\dagger}MC^{\dagger}CM^{*}B$. Conversely, suppose there exists some nonzero $w$ such that $B^{\dagger}MC^{\dagger}CM^{*}Bw=\lambda w$. Then $w\not\in\ker{B}$. With $v\coloneqq Bw$ it holds that $v\not=0$. By (36) and (39), $Tv=BB^{\dagger}MC^{\dagger}CM^{*}Bw=B\lambda w=\lambda v$. Thus $\lambda$ is an eigenvalue of $T$. We conclude that
@@ 4 Unbounded solution
Let $\mathcal{H}$ be a separable Hilbert space and take $\mathcal{H}_{i}\coloneqq\mathcal{H}$ for $1\leq i\leq 4$. Let $(e_{n})_{n}$ be an orthonormal basis of $\mathcal{H}$. Let $\0(\gamma_{n})_{n},(\mu_{n})_{n}\in\mathbb{C}\backslash\{0\}$ converge to zero. We assume $(\mu_{n})_{n}$ is nonincreasing. Let $(\alpha_{n})_{n}\in\mathbb{C}$ be such that $\lvert\alpha_{n}\rvert\rightarrow\infty$ and $(\alpha_{n}\gamma_{n})_{n}$ is square summable. For example, $\gamma_{n}=n^{-2}$ and $\alpha_{n}=n$ satisfy these conditions.
Define $w\coloneqq\sum_{n\geq 2}\alpha_{n}\gamma_{n}e_{n}\in\mathcal{H}$. Put $f_{1}\coloneqq w\lVert w\rVert^{-1}$ and $f_{2}\coloneqq e_{1}$ and extend $\{f_{1},f_{2}\}$ to an ONB $(f_{n})_{n}$ of $\mathcal{H}$. We then set, for $h\in\mathcal{H}$,
@@
Then $C$ is injective and $\operatorname{ran}{C}=\{h\in\mathcal{H}:\ \sum_{n}\gamma_{n}^{2}\langle h,e_{n}\rangle^{2}<\infty\}$. As $\operatorname{ran}{C}$ contains $\{e_{n}\}_{n}$, it is dense in $\mathcal{H}$, thus $\operatorname{ran}{C}^{\perp}=\{0\}$. Therefore, $\operatorname{dom}{C^{\dagger}}=\operatorname{ran}{C}$ and the Moore–Penrose inverse of $C$ is given, as in Example A.6, by
@@
Let us take $B=I$ for simplicity, so $P_{\overline{\operatorname{ran}{B}}}=I$. As $C$ is injective, $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}=M$. Since $(\mu_{n})_{n}$ is nonincreasing, for any $r\geq 1$, one of the rank-$r$ truncated SVDs of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ is given by
@@
$\geq\big{\lvert}\mu_{1}\gamma^{-1}_{m}\langle e_{m},f_{1}\rangle\big{\rvert}^{2}=\lvert\mu_{1}\rvert^{2}\lvert\alpha_{m}\rvert^{2}\lVert w\rVert^{-2}\rightarrow\infty$
as $m\rightarrow\infty$, showing that $B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C^{\dagger}$ is unbounded. This shows that (11) can indeed give unbounded solutions. If $\mu_{1}>\mu_{2}$ and $r=1$, then (11) gives a unique solution, which shows that (11) can also give only unbounded solutions. In this case, $\mathcal{S}_{r,*}$ only contains unbounded solutions, so that in fact all solutions of (10) are unbounded by Corollary 3.10.
Now let $r=1$ and $\mu_{1}=\mu_{2}$, so that $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ has at least two distinct rank-$1$ truncated SVDs. The rank-$1$ truncated SVD (18), i.e. $(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{1}=\mu_{1}\langle\cdot,f_{1}\rangle f_{1}$, yields the following solution, which is of the form (11):
@@
We have thus constructed two solutions of the form (11): one of these is bounded, in particular it has finite rank and has finite Hilbert–Schmidt norm; the other is unbounded.